Mr. Kim is 25 years older than his daughter. In 10 years Mr. Kim will be twice as old as his daughter will be. How old is each now?
Answer by macston(5194)

.
D=daughter's age; F=father's age=D+25
.
F+10yrs=2(D+10yrs) Substitute for F.
D+25yrs+10yrs=2D+20 yrs
D+35yrs=2D+20yrs
15yrs=D ANSWER 1: Daughter is 15 years old.
.
F=D+25yrs=15yrs+25yrs=40yrs ANSWER 2: Father is 40 years old
.
CHECK:
F+10yrs=2(D+10yrs)
40yrs+10yrs=2(15yrs+10yrs)
50yrs=2(26yrs)
50yrs=50yrs